OpenStudy (anonymous):

Medal! Just verify my answer please. If the square below has a perimeter of 32 in, find the area of the inscribed circle. answer = approximately 50.27in^2

OpenStudy (anonymous):



OpenStudy (jdoe0001):

does a square have 4 equal sides? or its sides are not equal to each other?

OpenStudy (anonymous):

 So radius = 4 Area = pi * r^2 = about pi1 * 4^2 = your answer